OpenStudy (jaydelv):

The point (−3, 1) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (anonymous):

Hm

OpenStudy (anonymous):

Hm

OpenStudy (jaydelv):

Im confused

OpenStudy (anonymous):

Same hold on i've got someone assisting me

OpenStudy (jaydelv):

Thank you (:

OpenStudy (anonymous):

tan theta=1/3

OpenStudy (jaydelv):

My internet just got back up

OpenStudy (jaydelv):

i don't understand how he did this

OpenStudy (jhannybean):

A terminal angle 

OpenStudy (jhannybean):

So now we just use SOHCAHTOA to find sin: cos: tan:

OpenStudy (jhannybean):

\[\sf \text{ sin }\theta = \frac{opp}{hyp} \\ \\ cos ~\theta = \frac{adj}{hyp} \\ \\ tan ~\theta =\frac{opp}{adj} \]

OpenStudy (anonymous):

use the Pythagorean theorem to find the radius, r²= x² + y²

OpenStudy (anonymous):

so now you just gotta plug in x and y from your point into the above formula